OpenStudy (anonymous):

Select the best possible first step to solving the system by first eliminating the x variable. 3x + 5y = 1 2x + 4y = –4

OpenStudy (anonymous):

3x=1-5y x=1-5y/3 2(1-5y/3)+4y=-4 6(1-5y)+4y=-4 6-30y+4y=-4

OpenStudy (anonymous):

-30y+4y=2 -26y=2 y=-1/13